A teaspoon of milk of magnesia contains 25.0 mg of magnesium hydroxide, Mg(OH)2.  What volume of 0.010 M HCl in a person’s stomach would be neutralized by this teaspoon of antacid?




1
Expert's answer
2021-01-22T06:22:49-0500

Solution:

The balanced chemical equation:

Mg(OH)2 + 2HCl = MgCl2 + 2H2O

According to the equation: n(Mg(OH)2) = n(HCl)/2


Mass of Mg(OH)2 = 25.0 mg = 0.025 g

Moles of Mg(OH)2 = Mass of Mg(OH)2 / Molar mass of Mg(OH)2

The molar mass of Mg(OH)2 is 58.32 g mol-1.

Hence,

n(Mg(OH)2) = 0.025 g / 58.32 g mol-1 = 0.0004287 mol


n(HCl) = 2 × n(Mg(OH)2) = 2 × 0.0004287 mol = 0.0008574 mol


Molarity of HCl = Moles of HCl / Volume of HCl

Hence,

Volume of HCl = Moles of HCl / Molarity of HCl

V(HCl) = 0.0008574 mol / 0.010 M = 0.08574 L = 85.74 mL

V(HCl) = 85.74 mL


Answer: 85.74 mL of HCl would be neutralized.